Liquid sodium is being considered as an engine coolant. How many grams of liquid sodium (minimum) are needed to absorb 9.10 MJ of energy in the form of heat if the temperature of the sodium is not to increase by more than 10.0 °C? Use ?P=30.8 J/(K⋅mol) for Na(l) at 500 K.
Answer #1

Given:
Q = 9100000 J
C = 30.8 J/mol.K = 30.8 J/mol.oC
ΔT = 10 oC
use:
Q = n*C*ΔT
9100000 = n*30.8*10.0
n = 2.95*10^4 mol


Molar mass of Na = 22.99 g/mol

use:
mass of Na,
m = number of mol * molar mass
= 2.95*10^4 mol * 22.99 g/mol
= 6.782*10^5 g
Answer: 6.78*10^5 g